OpenStudy (anonymous):

If S is the sum of integers from 51 to 150 and T is the sum of integers from 151 to 250, what is the value of S-T?

OpenStudy (whpalmer4):

Observe that 151 -51 = 100 152 - 52 = 100 ... 250 - 150 = 100

OpenStudy (anonymous):

okay

OpenStudy (whpalmer4):

though probably you should look at it like this: 51-151 = -100 52-152 = -100 ...

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

So, what's your answer?

OpenStudy (anonymous):

-10,000 by using formula

OpenStudy (whpalmer4):

Very good.